# RTLCoder: Fully Open-Source and Efficient LLM-Assisted RTL Code Generation Technique

### v4

@@ Abstract
The automatic generation of RTL code (e.g., Verilog) using natural language instructions and large language models (LLMs) has attracted significant research interest recently. However, most existing approaches heavily rely on commercial LLMs such as ChatGPT, while open-source LLMs tailored for this specific design generation task exhibit notably inferior performance. The absence of high-quality open-source solutions restricts the flexibility and data privacy of this emerging technique. In this study, we present a new customized LLM solution with a modest parameter count of only 7B, achieving better performance than GPT-3.5 on all representative benchmarks for RTL code generation. Especially, it outperforms GPT-4 in VerilogEval Machine benchmark. This remarkable balance between accuracy and efficiency is made possible by leveraging our new RTL code dataset and a customized LLM algorithm, both of which have been made fully open-source. Furthermore, we have successfully quantized our LLM to 4-bit with a total size of 4GB, enabling it to function on a single laptop with only slight performance degradation. This efficiency allows the RTL generator to serve as a local assistant for engineers, ensuring all design privacy concerns are addressed.
Manuscript received XXXX; revised XXXX; accepted XXXX. Date of publication XXXX; date of current version XXXX. This work is partially supported by Hong Kong Research Grants Council (RGC) ECS Grant 26208723, National Natural Science Foundation of China (62304192, 92364102), and ACCESS – AI Chip Center for Emerging Smart Systems, sponsored by InnoHK funding, Hong Kong SAR. (Corresponding author: Zhiyao Xie.)  Shang Liu, Wenji Fang, Yao Lu, Jing Wang, Qijun Zhang, Hongce Zhang and Zhiyao Xie are with the Department of Electronic and Computer Engineering at the Hong Kong University of Science and Technology (HKUST), Hong Kong SAR, China (email: sliudx@connect.ust.hk; wfang838@connect.ust.hk; yludf@connect.ust.hk; jwangjw@connect.ust.hk; qzhangcs@connect.ust.hk; hongcezh@ust.hk; eezhiyao@ust.hk).   Hongce Zhang is also with the Microelectronics Thrust at the Hong Kong University of Science and Technology (GZ), Guangzhou, China.
@@ I Introduction
In recent years, large language models (LLMs) such as GPT [1] have demonstrated remarkable performance in natural language processing (NLP). Inspired by this progress, researchers have also started exploring the adoption of LLMs in agile hardware design. Many new LLM-based techniques emerge and attract wide attention in 2023. For example, LLM-based solutions are proposed to generate design flow scripts to control EDA tools [2, 3], design AI accelerator architectures [4, 5], design quantum architectures [6], hardware security assertion generation [7], fix security bugs [8], and even directly generate the target design RTL [9, 10, 11, 12, 13, 14, 15, 3, 16, 17].
@@
Thakur et al. [13]
MG-Verilog [16] Open-Source Open-Source No
Goh et al. [17]
@@
Among the above explorations, a promising direction that perhaps attracts the most attention is automatically generating design RTL based on natural language instructions [9, 10, 11, 12, 13, 14, 15, 3, 16, 17]. Specifically, given design functionality descriptions in natural language, LLM can directly generate corresponding hardware description language (HDL) code [^1] Most existing works focus on generating design RTL in Verilog code. In this work, we also choose Verilog, while the method should be general and applicable to other HDL types like VHDL. We will use terms RTL code and Verilog code interchangeably. such as Verilog, VHDL, and Chisel from scratch. Compared with well-explored predictive machine learning (ML)-based solutions in EDA [20], such generative methods benefit the hardware design and optimization process more directly. This LLM-based design generation technique can potentially revolutionize the existing HDL-based VLSI design process, relieving designers from the tedious HDL coding tasks.
@@
Table I summarizes existing works in LLM-based design RTL generation. Some works [9, 10, 11, 14, 15] focus on prompt engineering methods based on commercial LLMs like GPT, without proposing new datasets or models for RTL code generation. As we will discuss later, reliance on commercial LLM tools limits in-depth research exploration and incurs serious privacy concerns in industrial IC design scenarios. Thakur et al. [13] generate a large unsupervised training [^2] Most customized LLM solutions (including RTLCoder) are developed by fine-tuning pre-trained LLMs based on a training dataset about the specific task. In this paper, we use the terms training and fine-tuning interchangeably. dataset by collecting Verilog-based projects from online resources like GitHub, then fine-tuning its own model. However, this unsupervised dataset is quite unorganized with a mixture of code and text. Evaluations on a third-party benchmark [11] show that the performance of its fine-tuned model is still inferior to commercial tools like GPT-3.5. The VerilogEval [12] from the NVIDIA research team proposes its own labeled training dataset and benchmark, then fine-tunes its own new model. This may be the first non-commercial model that claims comparable performance with GPT-3.5, but according to their authors, neither the training dataset nor fine-tuned LLM model will be released to the public in the near future [12]. Goh et al. [17] proposed an open-source dataset with Verilog instruction-code samples. However, their instructions have only module name information, without code functionality descriptions. A recently released dataset MG-Verilog [16] proposed a multi-grained dataset consisting of descriptions of various level details with code to enhance the model’s instruction-following ability. But the model performance of [16, 17] also cannot outperform GPT3.5. Besides these customized RTL-generation solutions, according to our study, all other software code (e.g., Python) generation models like CodeGen2 [21], StarCoder [22], and Mistral [23] are significantly inferior to GPT-3.5 in this RTL generation task.
@@
Compared with solutions based on closed-source commercial LLM tools like GPT, the open-source LLM solution is vitally important from both research and application perspectives: 1) For research purposes, obviously, closed-source commercial tools prevent most in-depth studies and customizations of this emerging technique. 2) For realistic applications, users of commercial LLM tools unavoidably have data privacy concerns, since all instructions have to be uploaded to LLM providers like OpenAI. The privacy concern is especially critical in the VLSI design industry, where information leakage of intellectual property (IP) or key technical innovations can seriously hurt the competitive advantage of users’ companies. In comparison, each user’s own local LLM developed based on an open-source solution can eliminate all privacy concerns and also ensure a reliable service.
@@
In this work, we finally fill this gap with our new open-source LLM solution named RTLCoder [^3] Our RTLCoder has been fully open-source at https://github.com/hkust-zhiyao/RTL-Coder. It includes the data generation flow, the complete generated training data set, the model training flow, and the final fine-tuned models (based on both Mistral and DeepSeek).. To the best of our knowledge, it is the first non-commercial and open-source LLM method that clearly outperforms GPT-3.5 in design RTL code generation. We validate this on two representative benchmarks [12, 11] and observe consistent trends. To build this RTLCoder, we first propose an automated data generation flow and have generated a dataset with over 27,000 instruction-code samples for the RTL generation task.
RTLCoder obviously achieves state-of-the-art trade-offs between performance and efficiency. Besides demonstrating unprecedented RTL generation correctness in non-commercial solutions, it only has 7 billion (B) parameters and can be trained with only 4 consumer-level GPU cards. After further quantizing the parameters to 4 bits, the RTLCoder-4bit takes only 4GB of memory and can work on a laptop with limited accuracy loss. As a result, our open-source lightweight RTLCoder solution is accessible to almost every research group and it is friendly to be implemented and improved by researchers. The contributions of RTLCoder can be summarized below:
  • Targeting Verilog code generation, we propose an automated flow to generate a large dataset with over 27 thousand diverse Verilog design problems and answers. It addresses the serious data availability challenge in IC design-related tasks, and its potential applications are not limited to LLMs. The LLM directly trained on it can already achieve comparable accuracy with GPT-3.5.
  • We introduce a new LLM training scheme based on code quality feedback. It further boosts the ultimate model performance to outperform GPT-3.5, being comparable with GPT-4. We further revised the training process from an algorithm perspective to reduce its GPU memory consumption. The training process only requires 4 commercial-level GPU cards.
  • We designed RTLCoder to be a lightweight solution with only 7B parameters. After quantizating its parameters into 4 bits, it takes only 4GB of memory, allowing it to serve as a local assistant for engineers without privacy concerns.
@@
Figure 1: Our proposed automated training dataset generation flow.
@@ II Automatic Dateset Generation
In this work, we first propose a new automated training dataset generation flow. Based on this flow, we have generated over 27 thousand training samples, with each sample being a pair of design description instruction (i.e., model input) and the corresponding reference RTL code (i.e., expected model output). The instruction can be viewed as the input question for LLMs, describing the desired circuit functionality in natural language. The reference code is the expected answer from LLMs, implementing the circuit functionality in Verilog code. We observe that these generated training samples exhibit high diversity and complexity in the RTL-generation domain, encompassing a diverse spectrum of difficulty levels.
We build this automated generation flow by taking full advantage of the powerful general text generation ability of the commercial tool GPT. Please notice that GPT is only used for dataset generation in this work, and we adopt GPT-3.5 in this data generation task. The automated dataset generation flow is illustrated in Figure 1, which includes three stages: 1) RTL domain keywords preparation, 2) instruction generation, and 3) reference code generation. We designed several general prompt templates to control GPT generating the desired outputs in each stage.
@@ II-A Stage 1: Keywords Preparation
Specifically, in this process 1, to collect a comprehensive range of RTL design task topics, we utilize a tree-like structure with multiple branches to issue queries to GPT. We first prompt GPT at the root node to provide categories and examples of frequently used block keywords in RTL design as Figure 5 illustrated. The response from GPT has a tree structure that consists of some subfields as Figure 3 shows. With the response, we could use the categories and examples as branches to continue prompting GPT for more design keywords within each topic. For example, we can use scripts to ask GPT about more types of the block “multiplier”, it will return more specific design names such as “Booth multiplier, Wallace tree multiplier, etc.”. After this process, we obtain hundreds of keywords related to RTL design in the Keywords pool $\mathcal{L}_{key}$.
Besides using our introduced keywords generation flow, we believe manually collecting initial Verilog instructions from existing IP families is also feasible and can help to produce a comprehensive keyword pool, as long as users devote manual efforts to collecting the IP document resources and checking their coverage and diversity.
This obtained keyword pool with about 350 instances will be extended into around 5K instruction samples in process 2 and these instructions will be further mutated and combined with each other and those obtained from source code in stage 2 to generate more new descriptions.
Figure 2: An example of Prompt $P_{key}$ in 1555We use red text boundary to denote GPT input examples, and green text boundary to denote GPT output examples in this work. Please notice that some green GPT output in this data generation flow are instructions, which will be the input of LLMs.
Figure 3: A GPT response example to Prompt $P_{key}$ in 1
@@ II-B Stage 2: Instruction Generation
The second stage targets generating sufficient instructions based on the initial keywords and Verilog source code. At process 2, we extend existing keywords from $\mathcal{L}_{key}$ to complete design instructions. Specifically, we randomly sample one or two keywords from $\mathcal{L}_{key}$ each time, combined with prompts $P_{ext}$, and feed them into GPT. The output is a complete RTL design instruction.
@@
Process 2 and 3 help generate the initial design instruction pool $\mathcal{L}_{ins}$ based on our customized prompt $P_{ext}$. Two types of prompt $P_{ext}$ are proposed for processing $\mathcal{L}_{key}$ and $\mathcal{L}_{code}$, denoted as $P_{ext}^{key}$ and $P_{ext}^{code}$, respectively. As shown in Figure 4, our prompt $P_{ext}^{key}$ in process 2 adopts the few-shot prompting technique, which means we provide an example of the question (i.e., keyword) and answer (i.e., instruction) in the input prompt. Figure 5 shows an example of GPT’s corresponding response. As for the prompt $P_{ext}^{code}$ used in process 3, an example of prompt and the response of GPT are provided in Figure 6 and Figure 7. The prompt $P_{ext}^{code}$ asks GPT to convert the given Verilog code snippet to the corresponding description instruction.
For process 3, instructions obtained through code snippets from $\mathcal{L}_{code}$ focus more on describing detailed circuit behavior (such as the specific behavior in different cycles of a signal). On the other hand, the instructions obtained from the keyword pool $\mathcal{L}_{key}$ in process 2 tend to provide an overall summary of the circuit’s functionality in a high-level manner. For example, based on the keyword “traffic light” in the keyword pool, a series of traffic light circuit instructions can be generated. These overall descriptions can hardly be obtained solely through code snippets if the signal name is not related to something like ’red_light’ or ’yellow_light’. The instructions generated through keywords not only describe circuit behavior but also link to commonly used digital design terminologies and higher-level abstracts. Hence, both initial instruction generation methods are crucial and complementary. They describe circuit behavior from different levels of description (highly abstract and detailed to the behavior of each signal).
In the extension process of 2 and 3, the prompt for generating the instruction responses should be carefully designed. We provide Figure 6 for the reader’s reference. The instructions returned by GPT should contain detailed and unambiguous requirements to ensure that LLM can generate only one functionally unique Verilog code based on this instruction. What’s more, the design complexity should be limited so that the corresponding Verilog code can be implemented in hundreds of code lines.
Figure 4: An example of Prompt $P_{ext}^{key}$ in 2
Figure 5: A GPT response example (green) to Prompt $P_{ext}^{key}$ when given topic: Pulse width modulators (PWM) (red) in 2
Figure 6: An example of Prompt $P_{ext}^{code}$ in 3
Figure 7: A GPT response example to Prompt $P_{ext}^{code}$ in 3.
@@
After generating the initial instruction pool $\mathcal{L}_{ins}$ with hundreds of initial instructions, we will iteratively use mutation methods to significantly augment the scale and complexity of this pool [^6] The scale and complexity in this context describe the target circuit. Circuit scale can be reflected by the total number of logic gates in the circuit after implementation. Circuit complexity can be reflected in its functionality complexity, which may be reflected by the number of lines in its implementation. Such metrics are also used in prior benchmarking works [11]. As an example, in Figure 10, we illustrate how combining a binary counter and a comparator can result in a circuit with a larger area and more complex functionality.. At 4, we use $P_{mut}$ to apply two types of mutation operations on instructions sampled from the design instruction library $\mathcal{L}_{ins}$. The process 4 would check every new design instruction using a set of rules and only passed valid instructions are added to $\mathcal{L}_{ins}$. We cover more details of this iterative process below.
For the mutation operation in 4, we propose two types of prompts $P_{mut}^{s}$ and $P_{mut}^{c}$, performing single circuit variation and circuits combination, respectively.
  • The $P_{mut}^{s}$ focuses on changing the single circuit module by requiring different functionalities while keeping the same hardware type, or implementing the same functionality but using different hardware. Figure 8 is an example of $P_{mut}^{s}$ and Figure 9 is a GPT’s response example with given instructions.
  • The $P_{mut}^{c}$ focuses on combining two circuit designs together. They are combined by either adding control logic to select using one of the two functionalities (i.e., parallel combination) or feeding one output into the other’s input (i.e., serial combination). Figure 10 gives a sample of prompt $P_{mut}^{c}$ and Figure 11 is a GPT’s response example with given instructions.
Figure 8: An example of Prompt $P_{mut}^{s}$ in 4
Figure 9: A GPT response example (green) to Prompt $P_{mut}^{s}$ with Given Instruction (red) in 4
It is important to note that we need to carefully design the instruction generation prompt considering the following two aspects in the mutation process:
  • Basic instruction validity777“Validity” here refers that the instruction can be implemented solely by Verilog code, requiring no other tools or languages.. For example, GPT may provide an instruction of implementing a physical temperature regulator or a speedometer, which are not directly related to Verilog coding.
  • Level of circuit behavior detail for LLM to generate correct code. Other studies [13, 16] have also indicated that the level of circuit description has a significant impact on the code generation quality. When descriptions are overly vague, LLMs trained on this dataset struggle to align the functionality of the code with the general instruction. Conversely, if descriptions are too detailed, focusing on intricate RTL circuit specifics, the RTL generation effectively becomes a form of “code translation”, which can also not boost the general generative abilities of trained LLMs.
Moreover, we request GPT to generate its reasoning steps (i.e., how it analyzes the code generation task step-by-step). These reasoning steps further enhance the detailed information of our instruction pool. In addition to incorporating these principles in the query prompt, we also use the one-shot technique (provide a query-response sample for GPT’s reference) to demonstrate our requirements. As long as following these principles and adopt basic prompt engineering techniques, we believe other alternative prompts may achieve similar response quality.
For the instruction checking in 5, we automatically check the correctness and diversity of each new instruction generated in the mutation process, based on a series of rules. The rules include 1) invalid-word checking, and 2) similarity measurement with existing $\mathcal{L}_{ins}$. For example, in invalid-word checking, if words such as “image” and “text” are detected, it indicates that GPT has likely provided a problem that cannot be implemented solely with RTL code. As for similarity rate checking, we calculate the similarity between the new instruction and existing $\mathcal{L}_{ins}$. The new instruction is discarded if the similarity is above a certain threshold since it can’t add enough new information to $\mathcal{L}_{ins}$. Such checking of instructions in 5 is imperfect, but it is fully automated and accurate enough to generate an ultimate instruction pool $\mathcal{L}_{ins}$, including over 50,000 instructions.
@@ II-C Stage 3: Reference Code Generation
Figure 10: An example of Prompt $P_{mut}^{c}$ in 4
Figure 11: A GPT response example (green) to Prompt $P_{mut}^{c}$ with Given Instruction (red) in 4
@@
The third stage targets generating the reference code. In the third stage, as shown in 6, we feed each instruction from $\mathcal{L}_{ins}$ into GPT, generating the corresponding reference design code as the solution. Specifically, we will ask GPT to provide at least five answers. After that, in 7, we will evaluate these answers using a code checker. In this work, we adopt an automated syntax checker to verify the correctness of RTL code syntax. Only syntax-correct design code can be kept. If all five answers fail the syntax checking, this instruction will be discarded. Finally, only valid instruction-code pairs are saved as our dataset. Ideally, process 7 should also check whether the functionality of the generated RTL code is consistent with the instruction, but currently generating testbenches for functionality verification cannot be automated. Similar to the checker in stage 2, this imperfect automated checking can already filter out the most serious mistakes in the dataset.
II-D Dataset Generation Flow Discussion
Figure 12: The circuit type distribution of keyword pool, the obtained dataset RTLCoder-27K, RTLLM-1.1 [11] and VerilogEval [12].
After going through all three proposed stages, we generate the ultimate training dataset named RTLCoder-27K with more than 27,000 instruction-code data samples. In this subsection, we will discuss some important topics about our proposed data generation flow and the output dataset. Multiple experiments are conducted to provide more insights about our RTLCoder dataset.
Statistics of keyword pool. In the keyword generation process, we adopt the hierarchical prompting structure to query GPT, resulting in around 350 keywords. Experienced engineers have recognized the keyword pool as covering a vast majority of commonly used circuit design problems. We also adopted GPT to analyze the keyword types and illustrate the categories and distribution of this keyword pool in Figure 12(a).
Comparison with other RTL dataset. We compare RTLCoder with other recent RTL code datasets [13, 17, 12, 16] in the following. In Thakur et al. [13], the training dataset only consists of code without instructions. Such code-only datasets do not align well with the RTL generation tasks based on natural language instructions. After removing exactly the same duplicated code in the dataset, there are approximately 25K data entries. In VerilogEval [12], each data entry comprises an instruction-code pair, totaling 8.5K entries in their dataset. However, their dataset is not open-sourced. In the MG-Verilog dataset [16], there are 11K instances with each one consisting of descriptions at various levels of detail and corresponding code. In the Goh et al. [17], there are a total of 60K samples. However, their instruction only provides the module name information without any functionality descriptions. The performance of LLMs [13, 12, 16, 17] trained on these 4 datasets will be presented in Table III in the following Section IV and all cannot clearly outperform GPT3.5. RTLCoder significantly outperforms all these models, indicating the overall quality of our open dataset.
Evaluation of design type distribution. We also utilized GPT for annotating the categories of the generated 27K-dataset. The circuit type distribution statistics is depicted in Figure 12(b). The distribution statistics for the two benchmarks, RTLLM-1.1 [11] and VerilogEval [12] are also plotted in Figure 12(c) and (d) respectively. The generated dataset consists of a wide range of circuit designs including Arithmetic Units, Signal Processing Units, Encryption Units, etc. It is important to note that a single data sample may implement functionalities of multiple circuit types, and we only adopt one circuit type to represent each instruction. As the majority of circuits contain submodules like Multiplexers, Decoders, and selectors, this may result in a relatively higher proportion of the Data Path type.
Evaluation of dataset diversity. To further check the diversity of our proposed training dataset, RTLCoder-27K, we utilized two diversity measures: Compression Ratios (CR) and Part-of-Speech Compression Ratio (CR: POS) which are suggested the best lexical diversity metrics by [25] among Homogenization Score (BERTScore), Self-BLEU, Homogenization Score (ROUGE-L), N-Gram Diversity Score, Hypergeometric Distribution, etc. CR is calculated utilizing text compression algorithms which can identify redundancy in the whole contents. We constructed a file $D$ containing all the text data in RTLCoder-27K following  [25]. Subsequently, we compressed the entire dataset using gzip. The Compression Ratio (CR) is as follows:
$\mathrm{CR}(D)=\frac{\text{ size of }D\oplus}{\text{ compressed size of }D\oplus}$
The CR-POS can capture the repeated syntactic redundancy by compressing the part-of-speech (POS) tag sequences of the original text. We also followed [25] and used NLTK POS tagger and the Penn Treebank set of 36 tags to extract the tag sequences [25].
The results are shown in Tabel II. Additionally, we also provide the CR and CR:POS values for three representative fine-tuning datasets: Goh et al. [17], MG-Verilog [16], and Magicoder-OSS-Instruct-75K [24] for comparison. A smaller CR and CR-POS value indicates a higher level of diversity in the dataset. We observe that RTLCoder-27K exhibits similar diversity to the widely used Python benchmark and higher diversity compared to the Verilog code datasets.
RTLCoder-
27K
Goh et al.
 [17]
MG-Verilog
 [16]
Magicoder-OSS-
Instruct-75K [24]
CR 4.41 5.27 5.80 4.02
CR: POS 7.61 10.1 9.16 6.67
Table II: The diversity scores of 4 representative fine-tuning instruct datasets. There are three Verilog datasets, including our RTLCoder-27K, Goh et al. [17] and MG-Verilog [16]. We also evaluate a widely used software (mainly Python) dataset, Magicoder-OSS-Instruct-75K [24]. The lower value of CR and CR: POS, the higher dataset diversity. RTLCoder-27K exhibits a satisfactory level of diversity compared with the other three datasets.
Imperfection in data functionality correctness. Our dataset generation flow ensures the syntax correctness of all instruction-code data samples, but cannot guarantee that every data sample is functionally correct (i.e., code implements the same functionality as described in instruction). This is because automatically checking the functionality correctness of training samples is inherently a very challenging task. Functionality checking for the Verilog code is practically hardware verification, which has been studied for decades, relies on human engineers, and is difficult to get guaranteed results. Although the code in our proposed dataset has only undergone syntax checks, finetuning the model using this dataset can still lead to improved model performance on the benchmarks as Table III shows in section IV. This indirectly demonstrates the contribution of the proposed dataset.
@@ II-D Dataset Generation Flow Discussion
Relation between RTLCoder and GPT-3.5. An interesting observation is that, although we generate our training dataset based on GPT-3.5, RTLCoder turns out to outperform the GPT-3.5 baseline on representative benchmarks [11, 12]. One important reason is that, for each instruction, we have employed a syntax checker to evaluate the reference code generated based on GPT-3.5. Therefore, among all correct and incorrect code from GPT-3.5, we filter out the obviously incorrect ones and retain the largely correct ones for training RTLCoder. This process can be viewed as a refinement of GPT-3.5’s Verilog generation capabilities.

### v5

# RTLCoder: Outperforming GPT-3.5 in Design RTL Generation with Our Open-Source Dataset and Lightweight Solution

@@ Abstract
The automatic generation of RTL code (e.g., Verilog) using natural language instructions and large language models (LLMs) has attracted significant research interest recently. However, most existing approaches heavily rely on commercial LLMs such as ChatGPT, while open-source LLMs tailored for this specific design generation task exhibit notably inferior performance. The absence of high-quality open-source solutions restricts the flexibility and data privacy of this emerging technique. In this study, we present a new customized LLM solution with a modest parameter count of only 7B, achieving better performance than GPT-3.5 on all representative benchmarks for RTL code generation. Especially, it outperforms GPT-4 in VerilogEval Machine benchmark. This remarkable balance between accuracy and efficiency is made possible by leveraging our new RTL code dataset and a customized LLM algorithm, both of which have been made fully open-source [^1] RTLCoder has been open-source in https://github.com/hkust-zhiyao/RTL-Coder. It includes the data generation flow, the complete generated training data set, the model training flow, and the final fine-tuned models (based on both Mistral and DeepSeek)..
publicationid: pubid: 979-8-3503-7608-1/24\$31.00 ©2024 IEEE \$*\$\$*\$footnotetext: Corresponding Author
@@ I Introduction
In recent years, large language models (LLMs) such as GPT [1] have demonstrated remarkable performance in natural language processing (NLP). Inspired by this progress, researchers have also started exploring the adoption of LLMs in agile hardware design. Many new LLM-based techniques emerge and attract wide attention in 2023. For example, LLM-based solutions are proposed to generate design flow scripts to control EDA tools [2, 3], design AI accelerator architectures [4, 5], design quantum architectures [6], hardware security assertion generation [7], fix security bugs [8], and even directly generate the target design RTL [9, 10, 11, 12, 13, 14, 15, 3].
Among the above explorations, a promising direction that perhaps attracts the most attention is automatically generating design RTL based on natural language instructions [9, 10, 11, 12, 13, 14, 15, 3]. Specifically, given design functionality descriptions in natural language, LLM can directly generate corresponding hardware description language (HDL) code such as Verilog, VHDL, and Chisel from scratch. Compared with well-explored predictive machine learning (ML)-based solutions in EDA [16], such generative methods benefit the hardware design and optimization process more directly. This LLM-based design generation technique can potentially revolutionize the existing HDL-based VLSI design process, relieving designers from the tedious HDL coding tasks.
@@
Thakur et al. [13] Open-Source Open-Source No
@@
Table I summarizes existing work on LLM-based design RTL generation. Some works [9, 10, 11, 14, 15] focus on prompt engineering methods based on commercial LLMs like GPT, without proposing new datasets or models for RTL code generation. As we will discuss later, reliance on commercial LLM tools limits in-depth research exploration and incurs serious privacy concerns in industrial IC design scenarios. Thakur et al. [13] generate a large unsupervised training [^2] Most customized LLM solutions (including RTLCoder) are developed by fine-tuning pre-trained LLMs based on a training dataset about the specific task. In this paper, we use the terms training and fine-tuning interchangeably. dataset by collecting Verilog-based projects from online resources like GitHub, then fine-tuning its own model. However, this unsupervised dataset is quite unorganized with a mixture of code and text. Evaluations on a third-party benchmark [11] show that the performance of its fine-tuned model is still inferior to commercial tools like GPT-3.5. The VerilogEval [12] from the NVIDIA research team proposes its own labeled training dataset and benchmark, then fine-tunes its own new model. This may be the first non-commercial model that claims comparable performance with GPT-3.5, but according to their authors, neither the training dataset nor fine-tuned LLM model will be released to the public in the near future [12]. Besides these customized RTL-generation solutions, according to our study, all other software code (e.g., Python) generation models like CodeGen2 [19], StarCoder [20], and Mistral [21] are significantly inferior to GPT-3.5 in this RTL generation task.
@@
Compared with solutions based on closed-source commercial LLM tools like GPT, the open-source LLM solution is vitally important from both research and application perspectives: 1) For research purposes, obviously, closed-source commercial tools prevent most in-depth studies and customizations of this emerging technique. 2) For realistic applications, users of commercial LLM tools unavoidably have data privacy concerns, since all instructions have to be uploaded to LLM providers like OpenAI. In comparison, each user’s own local LLM developed based on an open-source solution can eliminate all privacy concerns and also ensure a reliable service.
Figure 1: Our proposed automated training dataset generation flow.
@@
In this work, we finally fill this gap with our new open-source LLM solution named RTLCoder. To the best of our knowledge, it is the first open-source LLM that outperforms GPT-3.5 in all representative RTL code generation benchmarks [12, 11]. Our contributions are summarized below.
  • Targeting Verilog code generation, we propose an automated flow to generate a large labeled dataset with over 27 thousand diverse Verilog design problems and answers. It addresses the serious data availability challenge in IC design-related tasks, and its potential applications are not limited to LLMs. LLM directly trained on it can already achieve comparable accuracy to GPT-3.5.
  • We introduce a new memory-efficient LLM training scheme based on code quality feedback. It further boosts the ultimate model performance to outperform GPT-3.5, being comparable to GPT-4. Our 7B model can be trained with only four commercial GPU cards.
@@ II Automatic Dateset Generation
In this work, we first propose a new automated training dataset generation flow. Based on this flow, we have generated over 27 thousand training samples, with each sample being a pair of design instruction (i.e., model input) and the reference RTL code (i.e., expected model output). The instruction can be viewed as the input question for LLMs, describing the desired circuit functionality in natural language. The reference code is the expected answer from LLMs, implementing the circuit functionality in Verilog code. We observe that these generated training samples exhibit high diversity and complexity in the RTL-generation domain, encompassing a diverse spectrum of difficulty levels.
We build this automated generation flow by taking full advantage of the powerful general text generation ability of the commercial tool GPT. Please notice that GPT is only used for dataset generation in this work and GPT-3.5 is adopted here. The automated dataset generation flow is illustrated in Figure 1, which includes three stages: 1) RTL domain keywords preparation, 2) instruction generation, and 3) reference code generation. We designed several general prompt templates to control GPT generating the desired outputs in each stage.
@@ II-A Stage 1: Keywords Preparation
Specifically, in this process 1, to collect a comprehensive range of RTL design task topics, we utilize a tree-like structure with multiple branches to issue queries to GPT. We first prompt GPT at the root node to provide categories and examples of frequently used block keywords in RTL design. The response from GPT has a tree structure that consists of some related subfields. With the response, we could use the categories and examples as branches to continue prompting GPT for more design keywords within each topic. For example, we can use scripts to ask GPT about more types of the block “multiplier”, it will return more specific design names such as “Booth multiplier, Wallace tree multiplier, etc.”. After this process, we obtain hundreds of keywords related to RTL design in the Keywords pool $\mathcal{L}_{key}$.
@@ II-B Stage 2: Instruction Generation
The second stage targets generating sufficient instructions based on the initial keywords and Verilog source code. At process 2, we extend existing keywords from $\mathcal{L}_{key}$ to complete instructions. Specifically, we randomly sample one or two keywords from $\mathcal{L}_{key}$ each time, combined with prompts $P_{ext}$, and feed them into GPT to obtain an RTL design instruction.
@@
Process 2 and 3 help generate the initial design instruction pool $\mathcal{L}_{ins}$ based on our customized prompt $P_{ext}$. After generating the initial instruction pool $\mathcal{L}_{ins}$ with hundreds of initial instructions, we will iteratively use mutation methods to significantly augment the scale and complexity of this pool. At 4, we use $P_{mut}$ to apply two types of mutation operations on instructions sampled from the design instruction library $\mathcal{L}_{ins}$. The process 5 would check every new design instruction using a set of rules and only passed valid instructions are added to $\mathcal{L}_{ins}$. Stage 2 is fully automated and accurate enough to generate a high-quality ultimate instruction pool $\mathcal{L}_{ins}$, including over 50,000 instructions.
In addition, we will further request GPT to generate its reasoning steps (i.e., how it analyzes the generation task step-by-step). These reasoning steps further enhance the detailed information of our instruction pool.
@@ II-C Stage 3: Reference Code Generation
The third stage targets generating the reference code. In the third stage, as shown in 6, we feed each instruction from $\mathcal{L}_{ins}$ into GPT, generating 5 corresponding reference design codes as the solution candidates. After that, in 7, we will evaluate these answers using a code checker. In this work, we adopt an automated syntax checker and only syntax-correct design code can be kept. If all 5 answers fail the syntax checking, this instruction will be discarded. Finally, only valid instruction-code pairs are saved as our dataset. Ideally, process 7 should also check whether the functionality of the generated RTL code is consistent with the instruction, but currently generating testbenches for functionality verification cannot be automated. This imperfect automated checking can already filter out the most serious mistakes in the dataset.
@@
After going through all three stages, we generate the ultimate training dataset with more than 27,000 data samples. An interesting observation is that, although we generate our training dataset based on GPT-3.5, RTLCoder turns out to outperform GPT-3.5 on representative benchmarks [11, 12]. One important reason is that, for each instruction, we have employed a syntax checker to filter out the obviously incorrect codes generated from GPT-3.5 and retain the largely correct ones for training RTLCoder. This process can be viewed as a refinement of GPT-3.5’s Verilog generation capabilities.